OpenStudy (anonymous):

select the best answers, if f(x)=8x and g(x)=2x+1, find (f o g)(x) a. 10x+1 b. 16x^2+8x c. 16x+8 d. 16x+1

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(2x+1)=8(2x+1)=16x+8\]

OpenStudy (anonymous):

*answer

OpenStudy (anonymous):

therefor is it d?